3. Given the following reaction: Zn + CuCl2 → ZnCl2 + Cu

a. How many grams of ZnCl2 will be produced from 7.00 g of Zn and 8.00 g of CuCl2?

b. What is the limiting reagent?

c. What is the excess reagent and by how much?


1
Expert's answer
2021-12-14T00:14:02-0500

Molar mass of Zn=65.38

7/65.38=0.11

Molar mass of CUCl2=134.45 

8/134.45=0.06

Molar mass of ZnCl2= 136.286

=0.17×136.286= 23.17g